# EDiT: Efficient Diffusion Transformers with Linear Compressed Attention

### v1

@@ 4.1 Evaluations for EDiT
Hybrid CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ Latency (s) $\downarrow$
@@
SD-v3.5M 0.283 10.49 3.86 0.286 08.83 2.89 010.9
MM-EDiT (Ours) 0.285 11.60 3.91 0.287 09.73 2.95 08.00
MM-EDiT with $\eta^{Lin}$ 0.283 13.05 4.44 0.285 10.77 3.18 08.60
MM-EDiT no $\phi_{CF}$ and $\phi_{SC}$ 0.287 20.98 5.18 0.279 17.67 5.96 07.60
SANA-MM-DiT $\times$ 0.279 14.94 5.02 0.283 10.88 3.26 07.60
Linear MM-DiT-$\alpha$$\times$ 0.281 13.59 4.28 0.284 10.29 2.94 07.70
Linear MM-DiT-$\beta$$\times$ 0.281 13.53 4.37 0.281 09.75 2.94 07.75
Table 2: Quantitative comparisons of MM-EDiT to it’s teacher (SD-v3.5M) and alternative design choices. The latency indicates the time to generate a single $1024\times 1024$ pixel image on a consumer-grade Nvidia 3090 RTX with $28$ diffusion steps. Our results demonstrate that MM-EDiT performs at par with the teacher SD-v3.5M with significant improvement in latency.
@@
Figure 3: Qualitative results for EDiT and MM-EDiT models at a resolution of $1024\times 1024$ compared to base models Pixart-$\Sigma$ and SD-v3.5M, and SANA-based SANA-DiT and SANA-MMDiT respectively. The prompts used to generate these images and results for all other baselines and ablations can be found in Appendix A.
Baselines and Ablations. We compare the performance of our EDiT version distilled from PixArt-$\Sigma$ with the original PixArt-$\Sigma$ model. Additionally, we evaluate EDiT against Linfusion-DiT, which uses Linfusion’s attention mechanism [14] within PixArt-$\Sigma$, reproducing the results of [14]. We further consider SANA-DiT, which uses the linear attention and MixFFN from SANA [22] to augment the PixArt-$\Sigma$ architecture as well as the original KV Token Compression (KV Comp) approach from [2]. For consistency, we use our distillation approach for all architectures.
Finally, we evaluated EDiT’s design choices for computing queries, keys, and values using three settings: applying ConvFusion on queries and keys, using only Spatial Compressor on keys and values, and applying ConvFusion on queries without compressing keys and values. This helps us understand the impact of enhancing token interactions in the queries and shows that compressing does not harm performance.
Quantitative Results. Table 1 summarizes our EDiT results based on Pixart-$\Sigma$. At both $512\times 512$ and $1024\times 1024$ resolutions, EDiT performs on par with the teacher model (i.e., Pixart-$\Sigma$ baseline model) — exceeding its performance at the lower resolution. Additionally, our model surpasses the linear approaches SANA-DiT and Linfusion by 2.6 and 8.8 Inception-v3 FID points for $512\times 512$ resolution, respectively, and outperforms the KV Comp token compression method by 2.5 FID points. We observe a similar performance of EDiT on $1024\times 1024$ resolution. These results demonstrate that our EDiT model outperforms current linear attention and token compression approaches. Comparing EDiT to its ablations shows our mixture of ConvFusion and Spatial Compressor strikes an excellent balance between image quality and runtime. To show the effectiveness of our approach while generating high-resolution images, in Figure 4, we present the results for EDiT and PixArt-$\Sigma$ at a resolution of $2048\times 2048$, where EDiT performs comparably but requires around 2.5×2.5$\times$2.5 × less time.
PixArt-$\Sigma$
EDiT
CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ Runtime $\downarrow$
(Inception-v3) (Clip)
PixArt-$\Sigma$ 0.274 7.48 2.59 41.55
EDiT (ours) 0.288 8.29 2.71 16.85
Figure 4: Qualitative and quantitative comparision of EDiT and PixArt-$\Sigma$ at a resolution of $2048\times 2048$ pixels. EDiT generates comparable results whie achiving a speedup of almost 2.5×2.5$\times$2.5 ×.
4.2 Evaluations for MM-EDiT
Baselines and Ablations. Besides comparing our MM-EDiT-based SD-v3.5M with the original model, we compare MM-EDiT’s performance with several baselines and ablations. We compared MM-EDiT with three baselines using linear attention in MM-DiTs: Linear MM-DiT-$\alpha$, which computes queries, keys and tokens like MM-EDiT, but uses linear attention for all relations; Linear MM-DiT-$\beta$, which incorporates $\phi_{\textrm{LF}}$ from Linfusion for prompt tokens; and SANA-MM-DiT, which uses $\phi_{\textrm{SANA}}$ and MixFFN layers. Additionally, we considered two variations of our hybrid attention mechanism: MM-EDiT with $\eta^{Lin}$ (Eq 8), which uses theoretically correct adjustment terms, and MM-EDiT without $\phi_{CF}$ and $\phi_{SC}$. These comparisons highlight the importance of hybrid attention and token interaction in improving performance.
Quantitative Results Table 2 presents the performance of MM-EDiT based on SDv3.5M, comparing both hybrid and non-hybrid attention methods. Our approach outperforms non-hybrid methods such as SANA-MM-DiT, Linear MM-DiT-$\alpha$, and Linear MM-DiT-$\beta$, highlighting the importance of hybrid attention in MM-DiTs. Note that our Linear MM-DiT-$\beta$ approach achieves a comparable FID score, but the qualitative results are inferior than the MM-EDiT (see Appendix A). Among the hybrid techniques, using the true normalization factor $\eta^{Lin}$ yields slightly inferior performance and slower processing compared to our simple approximation. Additionally, when we replace $\phi_{CF}$ and $\phi_{SC}$ with basic linear functions, the performance drops considerably, highlighting the value of our convolution-based functions.
4.3 Qualitative Comparisons
@@
To demonstrate the visual quality of the images produced by EDiT and MM-EDiT, Figure 3 shows example images generated with our models. Appendix A provides the used prompts and images for the remaining approaches. These images demonstrate high quality, visually indistinguishable from that of the base model and high prompt adherence.
4.4 On-device Runtime Analysis
@@
In this section, we demonstrate the potential of our methods for reducing runtimes on edge devices. Table 5 presents a runtime analysis of MM-EDiT on a Qualcomm SM8750-AB Snapdragon 8 Elite—a processor used in state-of-the-art smartphones. We compare the runtimes of our EDiT and MM-EDiT with SANA-based baselines (SANA-DiT and SANA-MM-DiT) and their corresponding base models (Pixart-$\Sigma$ and SD-v3.5M). For DiTs, EDiT achieves a latency of 896 milliseconds—38% and 26% faster than base and SANA-DiT models, respectively. For MM-DiTs, MM-EDiT reduces the runtime of SD-v3.5M by $\times2$, although its runtime remains slightly higher than that of SANA-MM-DiT. Further investigation into MM-EDiT revealed that the scaled dot-product attention operations for both image-text $Q_{I}K_{P}^{T}$ and text-image $Q_{P}K_{I}^{T}$ interactions are computationally expensive on the device. This is primarily because token compression is not applied to image queries $Q_{I}$, making the dot-product calculation expensive. Even though text-image interactions $Q_{P}K_{I}^{T}$ are also resource-intensive, they incur a lower cost due to token compression applied to the image keys $K_{I}$. In contrast, SANA’s fully linear attention mechanism likely contributes to its slightly lower runtime compared to MM-EDiT, albeit at the expense of image quality. To address this, we encourage future research to explore alternative interaction strategies within hybrid attention for MM-DiTs.
@@ 5 Related Work
To the best of our knowledge CLEAR [13] is the only existing linear-time approach to MM-DiTs. However, they use sparse neighborhood attention [8], which requires intricate memory access patterns and often only results in practical efficiency gains for very high-resolution images ( $>2048\times 2048$ pixels). Further, it relies on specialized implementations, limiting its applicability to different hardware.
More broadly, linear DiTs can be combined with masking approaches to increase training efficiency as done for vanilla DiTs in PatchMixer [19], which is a promising research direction.
@@ 6 Conclusion
Limitations. Our evaluations were primarily conducted on PixArt-$\Sigma$ and Stable Diffusion 3.5-Medium. Future work should investigate the generalizability of our approach to larger models such as Flux [11]. Additionally, while quantitative metrics were used, the qualitative assessment of generated images remains subjective, and human evaluation could provide further insights into image quality.
Appendix A Prompts and Additional Qualitative Results
@@
    ”Super cool astronaut NFT, facing camera, stylish, bling glasses, photochromatic, plain background, futuristic 8k. >->- > Super cool astronaut NFT, facing camera, stylish, bling glasses, photochromatic, plain background.”

### v2

@@ 4.1 Evaluations for EDiT
Hybrid CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$
@@
SD-v3.5M 0.283 10.49 3.86 0.286 08.83 2.89
MM-EDiT (Ours) 0.285 11.60 3.91 0.287 09.73 2.95
MM-EDiT with $\eta^{Lin}$ 0.283 13.05 4.44 0.285 10.77 3.18
MM-EDiT no $\phi_{CF}$ and $\phi_{SC}$ 0.287 20.98 5.18 0.279 17.67 5.96
SANA-$M$-DiT $\times$ 0.279 14.94 5.02 0.283 10.88 3.26
Linear MM-DiT-$\alpha$ $\times$ 0.281 13.59 4.28 0.284 10.29 2.94
Linear MM-DiT-$\beta$ $\times$ 0.281 13.53 4.37 0.281 09.75 2.94
Table 2: Quantitative comparisons of MM-EDiT to it’s teacher (SD-v3.5M) and alternative design choices. Our results demonstrate that MM-EDiT performs at par with the teacher SD-v3.5M.
Baselines and Ablations. We compare the performance of our EDiT version distilled from PixArt-$\Sigma$ with the original PixArt-$\Sigma$ model. Additionally, we evaluate EDiT against Linfusion-DiT, which uses Linfusion’s attention mechanism [14] within PixArt-$\Sigma$, reproducing the results of [14]. We further consider SANA-DiT, which uses the linear attention and MixFFN from SANA [22] to augment the PixArt-$\Sigma$ architecture as well as the original KV Token Compression (KV Comp) approach from [2]. For consistency, we use our distillation approach for all architectures. Finally, we evaluated EDiT’s design choices for computing queries, keys, and values using three settings: applying ConvFusion on queries and keys, using only Spatial Compressor on keys and values, and applying ConvFusion on queries without compressing keys and values. This helps us understand the impact of enhancing token interactions in the queries and shows that compressing does not harm performance.
Quantitative Results. Table 1 summarizes our EDiT results based on Pixart-$\Sigma$. At both $512\times 512$ and $1024\times 1024$ resolutions, EDiT performs on par with the teacher model (i.e., Pixart-$\Sigma$ baseline model) — exceeding its performance at the lower resolution. Additionally, our model surpasses the linear approaches SANA-DiT and Linfusion by 2.6 and 8.8 Inception-v3 FID points for $512\times 512$ resolution, respectively, and outperforms the KV Comp token compression method by 2.5 FID points. We observe a similar performance of EDiT on $1024\times 1024$ resolution. These results demonstrate that our EDiT model outperforms current linear attention and token compression approaches. Comparing EDiT to its ablations shows our mixture of ConvFusion and Spatial Compressor strikes an excellent balance between image quality and runtime. To show the effectiveness of our approach while generating high-resolution images, in Figure 3, we present the results for EDiT and PixArt-$\Sigma$ at a resolution of $2048\times 2048$, where EDiT performs comparably but requires around $2.5\times$ less time.
PixArt-$\Sigma$
EDiT
CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$
(Inception-v3) (Clip)
PixArt-$\Sigma$ 0.274 7.48 2.59
EDiT (ours) 0.288 8.29 2.71
Figure 3: Qualitative and quantitative comparision of EDiT and PixArt-$\Sigma$ at a resolution of $2048\times 2048$ pixels. EDiT generates comparable results whie achiving a speedup of almost $2.5\times$.
4.2 Evaluations for MM-EDiT
Baselines and Ablations. Besides comparing our MM-EDiT-based SD-v3.5M with the original model, we also compare MM-EDiT’s performance with three other baselines using linear attention in MM-DiTs: Linear MM-DiT-$\alpha$, which computes queries, keys and tokens like MM-EDiT, but uses linear attention for all relations; Linear MM-DiT-$\beta$, which incorporates $\phi_{\textrm{LF}}$ from Linfusion for prompt tokens; and SANA-MM-DiT, which uses $\phi_{\textrm{SANA}}$ and MixFFN layers. Additionally, we considered two variations of our hybrid attention mechanism: MM-EDiT with $\eta^{Lin}$ (Eq 7), which uses theoretically correct adjustment terms, and MM-EDiT without $\phi_{CF}$ and $\phi_{SC}$. These comparisons highlight the importance of hybrid attention and token interaction.
Quantitative Results Table 2 presents the performance of MM-EDiT based on SDv3.5M, comparing both hybrid and non-hybrid attention methods. Our approach outperforms non-hybrid methods such as SANA-MM-DiT, Linear MM-DiT-$\alpha$, and Linear MM-DiT-$\beta$, highlighting the importance of hybrid attention in MM-DiTs. Note that our Linear MM-DiT-$\beta$ approach achieves a comparable FID score, but the qualitative results are inferior than the MM-EDiT (see Appendix D). Additionally, when we replace $\phi_{CF}$ and $\phi_{SC}$ with basic linear functions, the performance drops considerably, highlighting the value of our convolution-based functions.
4.3 Qualitative Comparisons
@@
Figure 4: Qualitative results for EDiT and MM-EDiT models at a resolution of $1024\times 1024$ compared to base models Pixart-$\Sigma$ and SD-v3.5M, and SANA-based SANA-DiT and SANA-MMDiT respectively. The prompts used to generate these images and results for all other baselines and ablations can be found in Appendix D.
@@
To demonstrate the visual quality of the images produced by EDiT and MM-EDiT, Figure 4 shows example images generated with our models. Appendix D provides the used prompts and images for the remaining approaches. These images demonstrate high quality, visually indistinguishable from that of the base model, and high prompt adherence.
$1024\times 1024$ $2048\times 2048$ $4096\times 4096$ $8192\times 8192$
PixArt-$\Sigma$ 0.047 0.387 4.770 72.96
SANA-DiT 0.043 0.166 0.687 21.76
EDiT (Ours) 0.034 0.121 0.461 1.693
SD-v3.5M 0.071 0.474 5.603
SANA-$M$-DiT 0.069 0.228 0.861
MM-EDiT (Ours) 0.053 0.169 0.623
Table 3: Latency measurements with varying image resolution on a NVIDIA A100. Here, we present the time taken for a single forward pass through the transformer component. Our results demonstrate that EDiT and MM-EDiT are significantly faster at higher resolutions. SD-v3.5M cannot generate $8K$ resolution images due to the limited size of positional embedding.
@@
4.4 Latency Analysis
Table 3 reports latency measurements for a single forward pass through the transformer module across baseline models and our proposed EDiT and MM-EDiT architectures, evaluated on a NVIDIA A100 GPU. We test a range of image resolutions, from $1024\times 1024$ up to $8192\times 8192$. At the highest resolution, our EDiT model achieves a latency of approximately 1.7 seconds—making it $43\times$ faster than Pixart-$\Sigma$ and $13\times$ faster than SANA-DiT. Similarly, our MM-EDiT approach is $9\times$ faster than SD-v3.5M when generating images at $4096\times 4096$ resolution. Notably, the latency improvement of EDiT and MM-EDiT over standard baselines scales quadratically with image resolution, underscoring their suitability for high-resolution image generation. Overall, EDiT and MM-EDiT strike the best trade-off between runtime and quality across all considered variants. Appendix B shows some additional runtime analysis on a consumer-grade GPU.
On-device Runtime Analysis In this section, we demonstrate the potential of our methods for reducing runtimes on edge devices. Figure 5 presents a runtime analysis of MM-EDiT on a Qualcomm SM8750-AB Snapdragon 8 Elite—a processor used in state-of-the-art smartphones. We compare the runtimes of our EDiT and MM-EDiT with SANA-based baselines (SANA-DiT and SANA-MM-DiT) and their corresponding base models (Pixart-$\Sigma$ and SD-v3.5M). EDiT achieves a latency of 896 milliseconds—38% and 26% faster than base and SANA-DiT models, respectively. MM-EDiT reduces the runtime of SD-v3.5M by $2.2\times$, although its runtime remains slightly higher than that of SANA-MM-DiT. Further investigation revealed that the scaled dot-product attention operations for both image-text $Q_{I}K_{P}^{T}$ and text-image $Q_{P}K_{I}^{T}$ interactions are computationally expensive on the device. This is primarily because token compression is not applied to image queries $Q_{I}$, making the dot-product calculation expensive. Even though text-image interactions $Q_{P}K_{I}^{T}$ are also resource-intensive, they incur a lower cost due to token compression applied to $K_{I}$. In contrast, SANA’s fully linear attention mechanism likely contributes to its lower runtime compared to MM-EDiT, albeit at the expense of image quality. To address this, we encourage future research to explore alternative interaction strategies within hybrid attention for MM-DiTs.
@@ 5 Related Work
To the best of our knowledge CLEAR [13] is the only existing linear-time approach to MM-DiTs. However, they use sparse neighborhood attention [8], which requires intricate memory access patterns and often only results in practical efficiency gains for very high-resolution images ( $2048\times 2048$ pixels). Further, it relies on specialized implementations, limiting its applicability to different hardware. More broadly, linear DiTs can be combined with masking approaches to increase training efficiency as done for vanilla DiTs in PatchMixer [19], which is a promising research direction.
@@ 6 Conclusion
Appendix A Limitations
Our evaluations were primarily conducted on PixArt-$\Sigma$ and Stable Diffusion 3.5-Medium. Future work should investigate the generalizability of our approach to larger models such as Flux [11]. Additionally, while quantitative metrics were used, the qualitative assessment of generated images remains subjective, and human evaluation could provide further insights into image quality.
Appendix B Additional Runtime Results
Model Latency (s)
PixArt-$\Sigma$ 5.44
EDiT (ours) 4.15
SANA-DiT 4.0
LinFusion-DiT 4.2
KV Comp. ($k=2$) 5.0
Ablations Q K V CF CF - - SC SC CF - -
4.55
3.8
4.3
Model Latency (s)
SD-v3.5M 10.9
MM-EDiT (Ours) 8.0
MM-EDiT with $\eta^{Lin}$ 8.6
MM-EDiT no $\phi_{CF}$ and $\phi_{SC}$ 7.6
SANA-$M$-DiT 7.6
Linear MM-DiT-$\alpha$ 7.7
Linear MM-DiT-$\beta$ 7.75
Table 4: Latency of generating a single $1024\times 1024$ pixel image on a consumer-grade Nvidia 3090 RTX. PixArt-$\Sigma$, EDiT, and all the respective ablations use 20 diffusion steps. SD-v3.5M, $M$-EDiT, and all the respective ablations use 28 diffusion steps. These results further emphasize that, while producing images of similar quality to their respective teachers, EDiT and MM-EDiT improve efficiency at the practically relevant resolution of $1024\times 1024$ and strike the best compromise of runtime and quality across all considered variants.
Table 4 shows additional runtime results for our models, the teachers, and all baselines and ablations for the relevant use-case of generating $1024\times 1024$ images on consumer-grade hardware.
Appendix C Finetuning
FID FID FID FID
(Inception-v3) CLIP (Inception-v3) CLIP
$512\times 512$ $1024\times 1024$
PixArt-$\Sigma$ 7.49 2.52 8.55 3.17
EDiT (before finetuing) 7.36 2.53 10.1 3.42
EDiT (after finetuing) 5.17 1.84 4.72 1.39
SD-v3.5 11.5 4.41 8.91 5.07
MM-EDiT (before finetuing) 8.43 2.81 11.6 5.28
MM-EDiT (after finetuing) 4.46 1.69 7.15 2.89
Table 5: Quantitative comparison of fine-tuning EDiT and MM-EDiT to fine-tuned pre-trained models. The results show that EDiT amd $M$-EDiT outperform fine-tuned baselines.
To show that our models are amenable to standard fine-tuning, we generate a high-quality synthetic dataset by generating $\approx 960,000$ images with Flux[Schnell] [11] re-using the YE-POP captions. From these, we use $\approx 24,000$ for validation and testing each. We then fine-tune our best distilled EDiT and MM-EDiT checkpoints using the remaining images while selecting the model using the validation set.
The results in Table 5 show that fine-tuning, as expected, improves FID on the test set, indicating that our models can not only be trained by disitillation but also using the standard diffusion losses of the teachers.
Appendix D Prompts and Additional Qualitative Results
@@
    ”Super cool astronaut NFT, facing camera, stylish, bling glasses, photochromatic, plain background, futuristic 8k. >-> Super cool astronaut NFT, facing camera, stylish, bling glasses, photochromatic, plain background.”